# Pseudo-Autoregressive Neural Codec Language Models for Efficient Zero-Shot Text-to-Speech Synthesis

### v1

@@ Abstract.
Recent zero-shot text-to-speech (TTS) systems face a common dilemma: autoregressive (AR) models suffer from slow generation and lack duration controllability, while non-autoregressive (NAR) models lack temporal modeling and typically require complex designs. In this paper, we introduce a novel pseudo-autoregressive (PAR) codec language modeling approach that unifies AR and NAR modeling. Combining explicit temporal modeling from AR with parallel generation from NAR, PAR generates dynamic-length spans at fixed time steps. Building on PAR, we propose PALLE, a two-stage TTS system that leverages PAR for initial generation followed by NAR refinement. In the first stage, PAR progressively generates speech tokens along the time dimension, with each step predicting all positions in parallel but only retaining the left-most span. In the second stage, low-confidence tokens are iteratively refined in parallel, leveraging the global contextual information. Experiments demonstrate that PALLE, trained on LibriTTS, outperforms state-of-the-art systems trained on large-scale data, including F5-TTS, E2-TTS, and MaskGCT, on the LibriSpeech test-clean set in terms of speech quality, speaker similarity, and intelligibility, while achieving up to ten times faster inference speed. Audio samples are available at https://anonymous-palle.github.io.
@@
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: arXiv Preprint; April, 2025; arXiv.orgccs: Computing methodologiesccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Natural language processingccs: Computing methodologies Natural language generation
Figure 1. Generation procedures of three paradigms, where generation proceeds top to bottom, and each row represents one step. (Left) AR generates tokens step by step from left to right until the [EOS] token. Grouped AR can predict multiple tokens per step. (Middle) NAR predicts all masked positions simultaneously in a temporally unordered manner, allowing partial token retention while remasking others for regeneration. (Right) PAR predicts all masked positions in parallel but commits only the leftmost span at each step, enforcing generation to follow the temporal order of speech.
@@ 1. Introduction
Text-to-speech (TTS) synthesis has made remarkable progress, driven by generative methods (Shen et al., 2018; Li et al., 2019; Kim et al., 2021; Ren et al., 2021; Jeong et al., 2021; Wang et al., 2023) and the expansion of available data (Guo et al., 2021; Ma et al., 2024; Chen et al., 2021; Yang et al., 2024c; Kang et al., 2024; He et al., 2024). These innovations have enabled TTS systems to achieve human-level parity in terms of naturalness and intelligibility, in both pre-defined specific speakers (Tan et al., 2024) and zero-shot scenarios (Chen et al., 2024a).
@@
Modern zero-shot TTS models can be broadly divided into two paradigms: autoregressive (AR) (Wang et al., 2023; Chen et al., 2024a; Meng et al., 2024; Wang et al., 2025; Yang et al., 2024a; Du et al., 2024a, c; Peng et al., 2024; Lajszczak et al., 2024; Anastassiou et al., 2024; Kharitonov et al., 2023; Kim et al., 2024; Liu et al., 2024; Sheng et al., 2025; Guo et al., 2024) [^1] For hybrid systems like VALL-E, we regard it as AR, as primary tokens are generated autoregressively. and non-autoregressive (NAR) (Shen et al., 2024; Ju et al., 2024; Le et al., 2023; Eskimez et al., 2024; Liu et al., 2025; Gao et al., 2023; Chen et al., 2024b; Wang et al., 2024; Casanova et al., 2022; Jiang et al., 2023, 2024, 2025; Lee et al., 2025). AR models sequentially generate speech from left to right, conditioning each step on previous outputs. Despite strong temporal grounding, AR models face fundamental limitations in generation efficiency and contextual modeling. Specifically, the fixed prediction size per step results in slow inference, with generation steps tightly coupled to the frame rate of output units. Additionally, the strictly unidirectional modeling restricts access to left context only and lacks future or global information, leading to exposure bias and posing challenges to robustness (Song et al., 2024a; Han et al., 2024; Du et al., 2024b). Techniques like grouped modeling (Chen et al., 2024a) and speculative decoding (Li et al., 2024) improve efficiency, but generation steps still scale linearly with synthesized speech duration. In contrast, NAR models enable fast inference through parallel generation. NAR TTS methods leverage denoising diffusion (Ju et al., 2024), flow matching (Eskimez et al., 2024; Chen et al., 2024b), GAN (Jiang et al., 2024), and masked generative modeling (Wang et al., 2024), drawing inspiration from advances in image (Chang et al., 2022) and language (Devlin et al., 2019) generation. Recent advances (Eskimez et al., 2024; Wang et al., 2024) demonstrate that text-to-speech alignment can be achieved without phoneme-level duration prediction or explicit alignment supervision. E2-TTS (Eskimez et al., 2024) and F5-TTS (Chen et al., 2024b) simply pad the character sequence with filler tokens to match the linearly predicted or user-specified speech length. MaskGCT (Wang et al., 2024) concatenates the tokenized text and speech sequences, adopts mask-and-predict training, and performs iterative parallel decoding during inference. However, despite their efficiency and simplicity, these methods lack temporal modeling and generate speech in a temporally unordered manner, resulting in slow convergence (Eskimez et al., 2024; Chen et al., 2024b), alignment inaccuracies (Eskimez et al., 2024), and degraded intelligibility (Wang et al., 2024).
@@
To combine the strengths of AR and NAR paradigms while mitigating their limitations, we propose a novel pseudo-autoregressive (PAR) codec language modeling paradigm that unifies AR and NAR modeling by combining temporal modeling from AR with parallel generation from NAR. PAR introduces a soft temporal inductive bias into an NAR structure, specifically a bidirectional masked generative transformer, by enforcing span-level causal ordering through progressive commitment. Local spans are generated simultaneously, with earlier ones committed before later ones, creating a span-level temporal flow that aligns with the structure of natural speech. Unlike grouped AR methods (Chen et al., 2024a), which generate fixed-length spans with dynamic time steps, PAR generates dynamic-length spans with fixed time steps. Building on PAR, we propose PALLE, a two-stage TTS system that performs PAR generation followed by confidence-guided refinement. In the first stage, PALLE progressively commits spans along the temporal dimension, generating a fixed portion of the target sequence at each step, resulting in constant inference steps regardless of target speech duration. In the second stage, low-confidence tokens are refined in parallel based on global context, requiring only a few steps to correct accumulated errors and yielding huge performance improvements.
@@
2. Related work
2.1. Zero-Shot TTS
Zero-shot TTS enables speech synthesis for unseen speakers by capturing the timbre, prosody, and style from short speech prompts. Benefiting from in-context learning (ICL), the performance of zero-shot TTS has improved significantly, reaching human-parity quality (Ju et al., 2024; Chen et al., 2024a). Recent approaches primarily fall into three categories: codec language modeling (Wang et al., 2023; Chen et al., 2024a; Wang et al., 2024), diffusion/flow-based methods (Shen et al., 2024; Ju et al., 2024; Le et al., 2023; Eskimez et al., 2024; Chen et al., 2024b), and hybrid approaches (Du et al., 2024a, c). Hybrid systems involve a text-to-codec language model with a codec-to-mel model based on diffusion or flow matching, disentangling complex speech components to improve zero-shot generalization.
@@ 2.2. Discrete Speech Representation
Discrete speech representations (Yang et al., 2024b) are typically categorized into acoustic tokens and semantic tokens. Acoustic tokens, produced by neural codec models (Défossez et al., 2023; Kumar et al., 2023; Zeghidour et al., 2022; Ji et al., 2025; Parker et al., 2024; Wu et al., 2024; Du et al., 2024d), aim to reconstruct speech with high fidelity, capturing fine-grained acoustic details. In contrast, semantic tokens are quantized from embeddings of self-supervised learning (SSL) models (Baevski et al., 2020a, b; Hsu et al., 2021; Chen et al., 2022), which are trained for discrimination or masked prediction, and preserve the linguistic content of speech. Recent methods have explored supervised semantic tokens (Du et al., 2024a, c), trained with automatic speech recognition (ASR) loss (Graves et al., 2006, 2013), making them more robust to noise and better suited for language modeling.
@@ 2.3. Neural Codec Language Modeling
Neural codec language modeling has emerged as a powerful paradigm for generative speech modeling. Built upon discrete speech representations, language modeling approaches are employed to predict speech tokens, which are then decoded into waveforms by a speech detokenizer. Two main paradigms have been explored: AR and NAR modeling. AR models (Wang et al., 2023; Chen et al., 2024a; Yang et al., 2024a; Du et al., 2024a, c; Kharitonov et al., 2023; Lajszczak et al., 2024; Sheng et al., 2025) generate tokens sequentially from left to right, providing strong temporal grounding but suffering from slow inference and limited context modeling. SyncSpeech (Sheng et al., 2025) is an AR masked generative model that incorporates temporal modeling via a causal attention mask and token-level duration prediction, aiming for streaming generation from partial text input. Both Grouped AR and SyncSpeech can predict multiple tokens at one time step, but the number of generation steps scales linearly with the input text, and only the previous context is considered during inference. In contrast, PAR generates the output sequence with fixed time steps while considering the global context. NAR masked generative models (Wang et al., 2024) improve generation efficiency by iteratively infilling masked positions using bidirectional context. MaskGCT (Wang et al., 2024) predicts the total speech length and performs temporally unordered prediction, relying solely on confidence-based scheduling, which may cause robustness issues in high-entropy regions such as fast transitions or expressive prosody.
@@ 3. Pseudo-Autoregressive Language Modeling
The PAR model is built upon a bidirectional masked generative transformer (Chang et al., 2022). During training, the model is optimized to maximize the likelihood of the speech tokens $\mathbf{c}$ conditioned on the text content $\mathbf{x}$ using cross-entropy loss. At each training step, a contiguous span mask is applied to $\mathbf{c}$ from a random starting position to the end:
@@
During inference, given a tokenized text $\mathbf{x}^{\text{gen}}$ to be synthesized, a speech prompt $\mathbf{y}^{\text{ref}}$ from an unseen speaker, and its corresponding tokenized text $\mathbf{x}^{\text{ref}}$, the speech tokens $\mathbf{c}^{\text{ref}}=[c^{\prime}_{0},c^{\prime}_{1},\ldots,c^{\prime}_{T^{\text{ref}}-1}]$ of length $T^{\text{ref}}$. The target speech token length $T^{\text{gen}}$ can be determined arbitrarily or estimated. Inference begins by initializing an extended speech token sequence:
@@
where the appended zeros serve as placeholders for the speech tokens to be generated. At each decoding step $t$, the model predicts all tokens in parallel, but retains only the leftmost $k^{\prime}=\min(\lfloor r^{\prime}T^{\text{gen}}\rfloor,N_{\text{left}})$ tokens, where $r^{\prime}\in(0,1)$ is a fixed ratio and $N_{\text{left}}$ is the number of tokens yet to be generated. Predictions are conditioned on the text prompt $\mathbf{x}^{\text{ref}}$, the target text $\mathbf{x}^{\text{gen}}$, and the current extended speech token sequence $\mathbf{c}^{\text{ext}}_{(t)}$. A binary mask is constructed as follows:
@@
Figure 3. Illustration of the two-stage PALLE framework. (a) The shared architecture: a bidirectional masked generative transformer. (b) Stage one: the model predicts all token positions in parallel but retains only the leftmost span at each step. (c) Stage two: the model refines the initial speech tokens, where low-confidence tokens are re-masked and regenerated using contextual information.
[…]
An illustration diagram of PALLE framework.
@@ 4.1. Architecture
Speech Tokenizer and Detokenizer We use the pretrained S3Tokenizer from CosyVoice 2 (Du et al., 2024a) to extract discrete semantic tokens from the input waveform at a rate of 25 Hz. This model is a fine-tuned version of the SenseVoice-Large (An et al., 2024) ASR model, trained on a large multilingual speech dataset for robust speech understanding. By leveraging ASR loss during training, the S3Tokenizer extracts semantic information, enabling implicit denoising and speaker disentanglement (Song et al., 2024b). We adopt the pretrained conditional flow matching (CFM) model from CosyVoice 2 (Du et al., 2024c) to reconstruct mel spectrograms from discrete speech tokens. The generated mel spectrogram is then converted into the waveform using the pretrained HiFi-GAN (Kong et al., 2020) vocoder from CosyVoice 2 (Du et al., 2024c).
@@ 4.3. Inference: In-Context Learning Via Prompting
Stage One The target speech token length $T^{\text{gen}}$ is either predefined or estimated linearly as in (Chen et al., 2024b):
@@
Table 1. Objective performance comparison on continuation and cross-sentence zero-shot speech synthesis tasks. Bold highlights the best result, while underlined marks the second-best. Metrics not reported in the original papers are calculated using the open-source checkpoints. GT duration denotes the results using the ground-truth total speech duration. For Continuation task, NAR systems without GT duration are excluded, as linear duration estimation requires forced alignment to obtain the transcription of the first three seconds of speech.
@@
VALL-E  (Chen et al., 2025) EnCodec (Défossez et al., 2023) Libriheavy - 3.00 0.770 - 3.80 0.630 0.73
@@
CosyVoice 2 (Sheng et al., 2025) S3Tokenizer v2 25Hz (Du et al., 2024c) LibriTTS - - - 3.00 - - 0.45
@@
Table 2. Subjective performance comparison on cross-sentence zero-shot speech synthesis tasks. GT duration denotes the results obtained using the ground-truth total speech duration.
@@
Ground Truth 4.23 0
E2 TTS (32 NFE) 3.95 -0.41
F5-TTS (32 NFE) 4.04 -0.20
MaskGCT 3.98 -0.25

### v2

@@ Abstract.
Recent zero-shot text-to-speech (TTS) systems face a common dilemma: autoregressive (AR) models suffer from slow generation and lack duration controllability, while non-autoregressive (NAR) models lack temporal modeling and typically require complex designs. In this paper, we introduce a novel pseudo-autoregressive (PAR) codec language modeling approach that unifies AR and NAR modeling. Combining explicit temporal modeling from AR with parallel generation from NAR, PAR generates dynamic-length spans at fixed time steps. Building on PAR, we propose PALLE, a two-stage TTS system that leverages PAR for initial generation followed by NAR refinement. In the first stage, PAR progressively generates speech tokens along the time dimension, with each step predicting all positions in parallel but only retaining the left-most span. In the second stage, low-confidence tokens are iteratively refined in parallel, leveraging the global contextual information. Experiments demonstrate that PALLE, trained on LibriTTS, outperforms state-of-the-art systems trained on large-scale data, including F5-TTS, E2-TTS, and MaskGCT, on the LibriSpeech test-clean set in terms of speech quality, speaker similarity, and intelligibility, while achieving up to ten times faster inference speed. Audio samples are available at https://microsoft.com/research/project/vall-e-x/palle.
@@
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia (MM ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3754745isbn: 979-8-4007-2035-2/2025/10ccs: Computing methodologiesccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Natural language processingccs: Computing methodologies Natural language generation
Figure 1. Generation procedures of AR, NAR, and PAR.
@@ 1. Introduction
Text-to-speech (TTS) synthesis has made remarkable progress, driven by generative methods (Shen et al., 2018; Li et al., 2019; Kim et al., 2021; Ren et al., 2021; Jeong et al., 2021; Wang et al., 2023) and the expansion of available data (Guo et al., 2021; Ma et al., 2024; Chen et al., 2021; Yang et al., 2025; Kang et al., 2024; He et al., 2024). These innovations have enabled TTS systems to achieve human-level parity in terms of naturalness and intelligibility, in both pre-defined specific speakers (Tan et al., 2024) and zero-shot scenarios (Chen et al., 2024).
@@
Modern zero-shot TTS models can be broadly divided into two paradigms: autoregressive (AR) (Wang et al., 2023; Chen et al., 2024; Meng et al., 2025; Wang et al., 2025a; Yang et al., 2024a; Sun et al., 2025; Wang et al., 2025b; Du et al., 2024a, b, 2025a; Peng et al., 2024; Lajszczak et al., 2024; Anastassiou et al., 2024; Kharitonov et al., 2023; Kim et al., 2024; Liu et al., 2024; Jia et al., 2025; Sheng et al., 2025; Guo et al., 2024; Deng et al., 2025; Zhou et al., 2025; Zhang et al., 2025; Ma et al., 2025) [^1] Hybrid systems like VALL-E are categorized by primary token generation. and non-autoregressive (NAR) (Shen et al., 2024; Ju et al., 2024; Le et al., 2023; Eskimez et al., 2024; Liu et al., 2025; Gao et al., 2023; Chen et al., 2025a; Wang et al., 2024; Casanova et al., 2022; Jiang et al., 2023, 2024, 2025; Lee et al., 2025; Zhu et al., 2025). AR models sequentially generate speech from left to right, conditioning each step on previous outputs. Despite strong temporal grounding, AR models face fundamental limitations in generation efficiency and contextual modeling. Specifically, the fixed prediction size per step results in slow inference, with generation steps tightly coupled to the frame rate of output units. Additionally, the strictly unidirectional modeling restricts access to left context only and lacks future or global information, leading to exposure bias and posing challenges to robustness (Song et al., 2025; Han et al., 2024; Du et al., 2025b). Techniques like grouped modeling (Chen et al., 2024) and speculative decoding (Li et al., 2025) improve efficiency, but generation steps still scale linearly with synthesized speech duration. In contrast, NAR models enable fast inference through parallel generation. NAR TTS methods leverage denoising diffusion (Ju et al., 2024), flow matching (Eskimez et al., 2024; Chen et al., 2025a), GAN (Jiang et al., 2024), and masked generative modeling (Wang et al., 2024), drawing inspiration from advances in image (Chang et al., 2022) and language (Devlin et al., 2019) generation. Recent advances (Eskimez et al., 2024; Wang et al., 2024) demonstrate that text-to-speech alignment can be achieved without phoneme-level duration prediction or explicit alignment supervision. E2-TTS (Eskimez et al., 2024) and F5-TTS (Chen et al., 2025a) simply pad the character sequence with filler tokens to match the linearly predicted or user-specified speech length. MaskGCT (Wang et al., 2024) concatenates the tokenized text and speech sequences, adopts mask-and-predict training, and performs iterative parallel decoding during inference. However, despite their efficiency and simplicity, these methods lack temporal modeling and generate speech in a temporally unordered manner, resulting in slow convergence (Eskimez et al., 2024; Chen et al., 2025a), alignment inaccuracies (Eskimez et al., 2024), and degraded intelligibility (Wang et al., 2024).
@@
To combine the strengths of AR and NAR paradigms while mitigating their limitations, we propose a novel pseudo-autoregressive (PAR) codec language modeling paradigm that unifies AR and NAR modeling by combining temporal modeling from AR with parallel generation from NAR. PAR introduces a soft temporal inductive bias into an NAR structure, specifically a bidirectional masked generative transformer, by enforcing span-level causal ordering through progressive commitment. Local spans are generated simultaneously, with earlier ones committed before later ones, creating a span-level temporal flow that aligns with the structure of natural speech. Unlike grouped AR methods (Chen et al., 2024), which generate fixed-length spans with dynamic time steps, PAR generates dynamic-length spans with fixed time steps. Building on PAR, we propose PALLE, a two-stage TTS system that performs PAR generation followed by confidence-guided refinement. In the first stage, PALLE progressively commits spans along the temporal dimension, generating a fixed portion of the target sequence at each step, resulting in constant inference steps regardless of target speech duration. In the second stage, low-confidence tokens are refined in parallel based on global context, requiring only a few steps to correct accumulated errors and yielding huge performance improvements.
@@
2. Related work
2.1. Zero-Shot TTS
Zero-shot TTS enables speech synthesis for unseen speakers by capturing the timbre, prosody, and style from short speech prompts. Benefiting from in-context learning (ICL), the performance of zero-shot TTS has improved significantly, reaching human-parity quality (Ju et al., 2024; Chen et al., 2024). Recent approaches primarily fall into three categories: codec language modeling (Wang et al., 2023; Chen et al., 2024; Wang et al., 2024), diffusion/flow-based methods (Shen et al., 2024; Ju et al., 2024; Le et al., 2023; Eskimez et al., 2024; Chen et al., 2025a), and hybrid approaches (Du et al., 2024a, b). Hybrid systems involve a text-to-codec language model with a codec-to-mel model based on diffusion or flow matching, disentangling complex speech components to improve zero-shot generalization.
@@ 2.2. Discrete Speech Representation
Discrete speech representations (Yang et al., 2024b) are typically categorized into acoustic tokens and semantic tokens. Acoustic tokens, produced by neural codec models (Défossez et al., 2023; Kumar et al., 2023; Zeghidour et al., 2022; Ji et al., 2025; Parker et al., 2024; Wu et al., 2024; Du et al., 2024c), aim to reconstruct speech with high fidelity, capturing fine-grained acoustic details. In contrast, semantic tokens are quantized from embeddings of self-supervised learning (SSL) models (Baevski et al., 2020a, b; Hsu et al., 2021; Chen et al., 2022), which are trained for discrimination or masked prediction, and preserve the linguistic content of speech. Recent methods have explored supervised semantic tokens (Du et al., 2024a, b), trained with automatic speech recognition (ASR) loss (Graves et al., 2006, 2013), making them more robust to noise and better suited for language modeling.
@@ 2.3. Neural Codec Language Modeling
Neural codec language modeling has emerged as a powerful paradigm for generative speech modeling. Built upon discrete speech representations, language modeling approaches are employed to predict speech tokens, which are then decoded into waveforms by a speech detokenizer. Two main paradigms have been explored: AR and NAR modeling. AR models (Wang et al., 2023; Chen et al., 2024; Yang et al., 2024a; Du et al., 2024a, b; Kharitonov et al., 2023; Lajszczak et al., 2024; Sheng et al., 2025) generate tokens sequentially from left to right, providing strong temporal grounding but suffering from slow inference and limited context modeling. SyncSpeech (Sheng et al., 2025) is an AR masked generative model that incorporates temporal modeling via a causal attention mask and token-level duration prediction, aiming for streaming generation from partial text input. Both Grouped AR and SyncSpeech can predict multiple tokens at one time step, but the number of generation steps scales linearly with the input text, and only the previous context is considered during inference. In contrast, PAR generates the output sequence with fixed time steps while considering the global context. NAR masked generative models (Wang et al., 2024) improve generation efficiency by iteratively infilling masked positions using bidirectional context. MaskGCT (Wang et al., 2024) predicts the total speech length and performs temporally unordered prediction, relying solely on confidence-based scheduling, which may cause robustness issues in high-entropy regions such as fast transitions or expressive prosody.
Figure 3. Illustration of the two-stage PALLE framework. (a) The shared architecture: a bidirectional masked generative transformer. (b) Stage one: the PAR model predicts all token positions in parallel but retains only the leftmost span at each step. (c) Stage two: the NAR model refines the initial speech tokens, where low-confidence tokens are re-masked and re-generated using contextual information.
[…]
An illustration diagram of PALLE framework.
@@ 3. Pseudo-Autoregressive Language Modeling
The PAR model is built upon a bidirectional masked generative transformer (Chang et al., 2022). During training, the model is optimized to maximize the likelihood of the speech tokens $\mathbf{c}$ conditioned on the text content $\mathbf{x}$ using cross-entropy loss. At each step, a contiguous span mask is applied to $\mathbf{c}$ from a random starting position to the end:
@@
During inference, given a tokenized text $\mathbf{x}^{\text{gen}}$ to be synthesized, a speech prompt $\mathbf{y}^{\text{ref}}$ from an unseen speaker, and its corresponding tokenized text $\mathbf{x}^{\text{ref}}$, the speech tokens $\mathbf{c}^{\text{ref}}=[c^{\prime}_{0},c^{\prime}_{1},\ldots,c^{\prime}_{T^{\text{ref}}-1}]$ of length $T^{\text{ref}}$ are extracted. The target speech token length $T^{\text{gen}}$ can be determined arbitrarily or estimated. Inference begins by initializing an extended speech token sequence:
@@
where the appended zeros serve as placeholders for the speech tokens to be generated. At each step $t$, the model predicts all tokens in parallel, but retains only the leftmost $k^{\prime}=\min(\lfloor r^{\prime}T^{\text{gen}}\rfloor,N_{\text{left}})$ tokens, where $r^{\prime}\in(0,1)$ is a fixed ratio and $N_{\text{left}}$ is the number of tokens yet to be generated. Predictions are conditioned on the text prompt $\mathbf{x}^{\text{ref}}$, the target text $\mathbf{x}^{\text{gen}}$, and the current extended speech token sequence $\mathbf{c}^{\text{ext}}_{(t)}$. A binary mask is constructed as follows:
@@ 4.1. Architecture
Speech Tokenizer and Detokenizer We use the pretrained S3Tokenizer from CosyVoice 2 (Du et al., 2024a) to extract discrete semantic tokens from the input waveform at a rate of 25 Hz. This model is a fine-tuned version of the SenseVoice-Large (An et al., 2024) ASR model, trained on a large multilingual speech dataset for robust speech understanding. By leveraging ASR loss during training, the S3Tokenizer extracts semantic information, enabling implicit denoising and speaker disentanglement (Song et al., 2024). We adopt the pretrained conditional flow matching (CFM) model from CosyVoice 2 (Du et al., 2024b) to reconstruct mel spectrograms from discrete speech tokens. The generated mel spectrogram is then converted into the waveform using the pretrained HiFi-GAN (Kong et al., 2020) vocoder from CosyVoice 2 (Du et al., 2024b).
@@ 4.3. Inference: In-Context Learning Via Prompting
Stage One The target speech token length $T^{\text{gen}}$ is either predefined or estimated linearly as in (Chen et al., 2025a):
@@
Table 1. Objective performance comparison on continuation and cross-sentence tasks. Bold highlights the best result, while underlined marks the second-best. Metrics not reported in the original papers are calculated using the open-source checkpoints. GT duration uses ground-truth total speech duration. For Continuation task, NAR systems without GT duration are excluded, as linear duration estimation requires forced alignment to obtain the transcription of the first three seconds of speech.
@@
VALL-E  (Chen et al., 2025b) EnCodec (Défossez et al., 2023) Libriheavy - 3.00 0.770 - 3.80 0.630 0.73
@@
CosyVoice 2 (Sheng et al., 2025) S3Tokenizer v2 25Hz (Du et al., 2024b) LibriTTS - - - 3.00 - - 0.45
@@
Table 2. Subjective performance comparison on cross-sentence zero-shot speech synthesis tasks. GT duration denotes the results obtained using the ground-truth total speech duration. CI (95%) shows 95% confidence intervals. p-value is obtained from the Wilcoxon signed-rank test to evaluate statistical significance.
@@
Ground Truth 4.23 0 0.18 -
E2 TTS (32 NFE) 3.95 -0.41 0.20 0.031
F5-TTS (32 NFE) 4.04 -0.20 0.18 0.033
MaskGCT 3.98 -0.25 0.17 0.025